Suppose that the demand curve facing OPEC is given by P=120-2Q and that each members coat of producing oil is AC=MC=$20. Find the cartels profit maximizing total output and price.

If instead of keeping to this output,all members overproduced their quotas by 20 percent,what would be the effect on OPEC's total profit?


1
Expert's answer
2021-12-05T18:59:30-0500

The cartels profit maximizing total output and price are:

MR = MC,

"MR = TR'(q) = 120 - 4q,"

120 - 4q = 20,

q = 25 units,

"P = 120 - 2\u00d725 = 70."

If instead of keeping to this output, all members overproduced their quotas by 20 percent, then the OPECS total profit will decrease.